PartnerObserving Log for 2007 Mar 17/18

Observers: Rick Pogge, Eichii Egami
Support: John Hill on-site, Olga Kuhn on-call
TO: Aaron Ceranski

Afternoon weather is warm with broken overcast, combination of thick high clouds and cumulus at midlevel building in place. Humidity 23%. Clouds are part of a system coming up from the SSW that began over Mexico this morning. Not looking good.

JMH: In the afternoon I increased the shim under the cryostat flange from 380 microns to 405 mincrons. This has the effect of raising the lower edge of chips 1-3 by ~10 microns.

All dates below are UTC 2007 March 18

0130 - Overcast above us, some virga rain hanging below the clouds to the W and NW of us, but humidity still ~30%. Pretty much a loss as sunset, might start to clear up around 3 or so during the Pluto occultation time. I doubt we'll get anything tonight during the SDT time (sunset to 2:30am).

0205 - 0400 - Mostly unsuccessful tests to get trackers and shutter to work together. These will serve as bias frames for Tech chip 1. We eventually succeed in getting the trackers to take 1 sec images on 4 sec centers. The main trick was to set the limits in /home/lbccontrol/guide.dat, but we also hacked the lbc program to not increase the tracker exposure time.

0430 - Totally overcast, wind ~4m/s with occasional gusts enough to rattle the building, humidity 40%.

0600 - Starting to see stars low in the S, but lots of high thick stuff covering most of the sky, and wind still gusty from the SSW.

0624 - Taking 25bias (might as well check out the camera while waiting for clouds to blow by).

0700 - Dry enough to open, but clouds still over most of the sky. Might as well equilibrate the mirror and get it ready in case it clears.

0710 - Running some tests of OBs to put a 4th mag star (a) on Chip 2, then (b) offset to technical chip 1, then (c) offset to another part of tech chip 1, (d) another to tweak the offset back (wrong way the other time)... a by-hand primary tweak of dZ4=2000 then another shot to see if it was right, and another dZ4=2000. This to setup for the pluto occultation later. The 4th mag star was barely visible in 60s through the clouds.

0736 - Winds are calm now, temperature still warm (7C), only 36% humidity, but overcast and pretty thick. A few bright stars barely discernable.

0750 - getting some stars, enough to adjust the primary with Vfastextra. We're way out, and there are still clouds, so this is a chance to learn more about how to "read" pupil images.

0853 - Done (some wait for clouds), got reasonable collimation. Tried to verfiy focus with Vfinefoc, but variable seeing clobbered it. Seeing ~1.2, but seeing worse on bottom of chip2, better on top. Dialing +500 of Z4... Made things worse, taking it out, it got better.

0910 - Going to M81, because it is still crappy and we have 30min to kill before handing off for the Pluto Occultation. Shooting at V. Long tweak up as I forgot to go to the offset position (d'oh!).
  • Image1: 092131 - Seeing 1.4", reasonably round & sharp stars - better than last night
  • Image2: 092308 - 1.4" w/round images
  • Image3: 092446 - 1.5" w/round images
  • Image4: 092624 - 1.5" w/round images
Seeing is worse than last night, but overally 2d image quality and uniformity is better.

0930 - It is now 0230MST, time to hand off to Hill & Babcock for the Pluto occultation.

-- RichardPogge - 17 Mar 2007

Pluto Occultation

Observers: Bryce Babcock and John Hill

Conditions are thin broken clouds and horrible seeing. (Herbst's record of 4.5 arcsec seeing may fall.)

0929 - Turn off camera; lbckill; Change bluechannel.conf, lbcnew, and guide.dat to force 1 sec guide exposures only; lbcstart; turn on.

0942 - Pluto_Rfastextra_Early to focus at an elevation similar to mid-occultation. 2 exposures. Did not send result to mirror since it was asking for 1500 nm Z7 on nice pupils.

0950 - Pluto_R3dither, 5 sec Sloan r' images of the Pluto field through clouds with horrible seeing

0957 - Pluto_RTech1, a long exposure which we throw away to create 1 sec guider images on Technical Chip 1. The interesting data is *_1.fits starting at 09:57 UT. Images are spaced at approximately 4 second intervals. We couldn't identify Pluto among all the other fuzzy blobs in the field so we offset to a brighter star in the field.

1029 - Star_RTech1_4k; Pluto now clearly identified in the field so we stay at this position for the occultation observation over the next 50 minutes.

1051 - Guider glitches and moves Pluto out of the tech chip frame. Lose ~3 minutes restarting the OB. Unfortunately this is very near the time of emmersion.

1054 - Star_RTech2_4k running again

1109 - Restart OB because LBC-Blue was apparently hung. In fact, it was still taking images but not transferring them. Lose ~3 minutes restarting the OB.

1124 - End of the 1 sec tech chip exposures.

1130 - Pluto_Rfastextra trying to focus on the Pluto field. Incredible numbers of overlapping pupils make it pretty hopeless.

1145 - Pluto_RVB1, 1 sec dithered images in r', V and B on the science chips.

1157 - Pluto_RVB1, 2 sec dithered images in r', V, B.

1215 - Camera RPC error. Turn off, lbckill, lbcstart, Turn on.

1230 - Sky_RVB1, 3 color dithered exposures of a field out of the Milky Way to use as flats.

1242 - Sky_Rtech1, 1 sec tracker exposures of the sky to simulate flats on the Tech chips.

1258 - End of the tracker flats at 15K counts/sec.

Close: Wind calm, sky clear, T=3.5degC.

1258 - 5 biases after restoring LBC-Blue software to its normal configuration.

-- JohnHill - 18 Mar 2007
Topic revision: r8 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback